River Enterprises has $503 million in debt and 23 million shares of equity outstanding. Its excess cash reserves are $14 millAlso,

2. If the growth rate is 3% the price per share is ?

3. If you are right and the growth rate is 3%, the price per share would be __$ higher?
Answer #1

If growth rate is 2% price per share=(196/(13%-2%)-503+14)/23=56.20948617

If growth rate is 3% price per share=(196/(13%-3%)-503+14)/23=63.95652174

Price would be higher by 63.95652174-56.20948617=7.74703557